# Visibility-Aware RRT* for Safety-Critical Navigation of Perception-Limited Robots in Unknown Environments

### v1

@@ IV GLOBAL PATH PLANNER
In this section, we introduce the Visibility-Aware RRT* algorithm, which aims to solve the motion planning problem under limited sensing capabilities while satisfying the visibility condition given in (16). The proposed algorithm builds upon the LQR-RRT* framework [13] by incorporating two CBFs that guide the search process to ensure satisfaction of both the collision-free (13) and visibility-aware (16) properties.
@@ IV-A Visibility-Aware RRT*
for $i=1,\ldots,\texttt{{maxIter}}$ do
@@
The LQR-RRT* [13] is a sampling-based motion planning algorithm that incorporates the LQR controller as a steering function for nonlinear systems with non-holonomic constraints. It guarantees probabilistic completeness and inherits the asymptotic optimality property from the RRT* algorithm through its rewiring process [14]. Recently, Yang et al. [15] proposed an extension to the LQR-RRT* that incorporates CBF constraints during the planning process. Instead of formulating and solving Quadratic Programs (QPs) iteratively, as done in existing works [16], it checks whether the CBF constraints are satisfied and uses them as the termination condition for the steering process.
  • Sample: Randomly samples a state ${\bm{x}}_{\text{rand}}=[x_{\text{rand}},y_{\text{rand}},\theta_{\text{rand}}]^{\top}$ within the configuration space to expand the tree towards.
  • ChooseParent: Selects the best parent node ${\bm{x}}_{\text{min}}$ for the new node ${\bm{x}}_{\text{new}}$ from the set of nearby nodes $\mathcal{X}_{\text{near}}$ based on the cost and the feasibility of the connecting trajectory generated by the steering function. Given a trajectory $\bm{\sigma}=({\bm{x}}(t),{\bm{u}}(t))$, where $t\in[0,T]$ and $T$ is the duration of the trajectory, the evaluated cost is defined as:
    $\texttt{cost}(\bm{\sigma})=\int_{0}^{T}\left({\bm{x}}(t)^{\top}Q{\bm{x}}(t)+{\bm{u}}(t)^{\top}R{\bm{u}}(t)\right)\,\mathrm{d}t,$ (17)
    where the $Q$ and $R$ are the same weight matrices as in (3) (see Alg. 2).
  • Rewire: Performs the rewiring process by considering the new node ${\bm{x}}_{\text{new}}$ as a potential parent for the nearby nodes in $\mathcal{X}_{\text{near}}$. If a trajectory from ${\bm{x}}_{\text{new}}$ to any nearby node ${\bm{x}}_{\text{near}}\in\mathcal{X}_{\text{near}}$ is found with a smaller cost (17), the tree is optimized by rewiring ${\bm{x}}_{\text{near}}$ to ${\bm{x}}_{\text{new}}$ (see Alg. 3).
@@
The LQR-CBF-Steer is the key function that allows the algorithm to use LQR (5) and CBF (6) as an extension heuristic between two nodes ${\bm{x}}_{\text{start}}$ and ${\bm{x}}_{\text{next}}$ (see Alg. 4). It first linearizes the dynamics around ${\bm{x}}_{\text{next}}$ (2) and computes the optimal gain $K_{\text{lqr}}$ using the Riccatti equation (5). The function then employs the LQR controller $\mathcal{K}_{\text{lqr}}^{*}$ to generate a sequence of intermediate states $\mathcal{X}^{\prime}=\{{\bm{x}}_{\text{start}},{\bm{x}}^{\prime}_{1},\ldots,{\bm{x}}^{\prime}_{T}\}$ and control inputs $\mathcal{U}^{\prime}=\{{\bm{u}}^{\prime}_{1},\ldots,{\bm{u}}^{\prime}_{T}\}$, steering the state from ${\bm{x}}_{\text{start}}$ towards ${\bm{x}}_{\text{next}}$ based on the linearized dynamics. Throughout this process, it verifies that all CBF constraints are satisfied for each state-action pair $({\bm{x}}^{\prime}_{t},{\bm{u}}^{\prime}_{t})$ at each discretized time step $t$. If any constraint is violated, the steering extension is immediately terminated [15].
@@ IV-B HOCBF for Collision Avoidance
In this work, we employ the unicycle model to demonstrate the proposed algorithm; however, the algorithm can be derived in the same way for more complex system dynamics that include rotational motion. The unicycle model is as follows:
@@
Hereafter, we abuse the notation and use ${\bm{x}}_{t}$ and ${\bm{u}}_{t}$ to refer to ${\bm{x}}^{\prime}_{t}$ and ${\bm{u}}^{\prime}_{t}$ in the context of the LQR-CBF-Steer function (Alg. 4). Let one of the obstacles with radius $r_{\text{obs}}\in\mathbb{R}^{+}$ be located at $(x_{\text{obs}},y_{\text{obs}})\in\mathbb{R}^{2}$. The constraint function (7a) for collision avoidance, considering the robot radius $r_{\text{r}}$ and the maximum tracking error $\epsilon$ mentioned in Sec. III, is defined as:
$h_{\textup{col}}({\bm{x}}_{t})=(x_{t}-x_{\text{obs}})^{2}+(y_{t}-y_{\text{obs}})^{2}-(r_{\text{obs}}+r_{\text{r}}+\epsilon)^{2}.$ (19)
Given the dynamics (18), the function $h_{\textup{col}}({\bm{x}}_{t})$ (19) has relative degree $r_{b}=1$ for the control input $v$ and $r_{b}=2$ for the control input $\omega$. To simplify the construction of the CBF constraint, we fix the translational velocity $v$ to a constant value and only control the angular velocity $\omega$ [16].
Then, the HOCBF constraint (9) with $r_{b}=2$ is [16]:
@@
$+k_{1}L_{f}h_{\textup{col}}({\bm{x}}_{t})+k_{2}h_{\textup{col}}({\bm{x}}_{t})\geq 0,$ (20)
where $k_{1}$ and $k_{2}$ are positive constants that are designed to satisfy the HOCBF condition (9), and
@@ IV-C1 Time-to-Reach
Given the robot’s current position ${\bm{r}}_{t}=[x_{t},y_{t}]^{\top}$ and the critical point ${\bm{x}}_{c}$, let $\Delta{\bm{r}}=\sqrt{(x_{t}-x_{c})^{2}+(y_{t}-y_{c})^{2}}$ denote the Euclidean distance between them. To account conservatively for the robot radius $r_{\text{r}}$ and the maximum tracking error $\epsilon$ of the local controller $\pi$, we define $\Delta d=\Delta{\bm{r}}-r_{\text{r}}-\epsilon$. The time-to-reach is:
@@ IV-C2 Time-to-Rotate
where $\bar{\omega}_{t}$ is the average angular velocity computed by simulating the LQR controller $\mathcal{K}_{\text{lqr}}^{*}$ for a rotation of $\Delta\theta$ starting from time $t$:
$\bar{\omega}_{t}=\frac{1}{T_{\omega}}\sum_{\tau=t}^{t+T_{\omega}}\omega_{\tau}^{\text{lqr}},$ (26)
with $\omega_{\tau}^{\text{lqr}}$ denoting the angular velocity generated by the LQR controller $\mathcal{K}_{\text{lqr}}^{*}$ at time step $\tau$, and $T_{\omega}$ being the total number of time steps required for the rotation.
@@ IV-C3 Visibility CBF
Then, the CBF constraint with $r_{b}=1$ (6) is defined as:
$\psi_{\textup{vis}}({\bm{x}}_{t})=\dot{h}_{\textup{vis}}({\bm{x}}_{t})+k_{3}\,h_{\textup{vis}}({\bm{x}}_{t})\geq 0,$ (28)
where $k_{3}$ is positive constant that is designed to satisfy the CBF condition. The gradient $\dot{h}_{\textup{vis}}({\bm{x}}_{t})$ can be derived as:
@@
To demonstrate that this is a valid CBF, let us verify that (6) holds for all ${\bm{x}}\in\mathcal{X}$. $\frac{x_{t}-x_{c}}{\Delta{\bm{r}}}$, $\frac{y_{t}-y_{c}}{\Delta{\bm{r}}}$, $\cos\theta_{t}$ and $\sin\theta_{t}$ range between $-1$ and $1$. Plus, $\left|\frac{\sin\theta_{t}\cos\theta_{c}-\cos\theta_{t}\sin\theta_{c}}{\sqrt{1-z_{t}^{2}}}\right|\leq 1$. $\bar{\omega}_{t}$ is bounded since it is computed from the average LQR control input, which converges to zero over time, and the largest control input is bounded in the control space $\mathcal{X}$. Provided that all terms in (28) are bounded and $h_{\textup{vis}}({\bm{x}}_{t})\geq 0$, we can choose $k_{3}$ such that the condition (6) is satisfied for all ${\bm{x}}_{t}\in\mathcal{X}$. As a result, $h_{\textup{vis}}$ is a valid CBF.
@@ V LOCAL TRACKING CONTROLLER
After the global planning cycle described in Sec. IV, the local controller $\pi$ tracks the waypoints ${\bm{x}}_{i}^{\text{ref}}\in\mathcal{P}$ enabling the robot to navigate towards the goal. During navigation, the robot utilizes its onboard sensor to create a map of the environment and detect obstacles on-the-fly. The environment may contain unknown obstacles that were not considered during the global planning phase. The local controller must ensure that the robot remains within the true collision-free set $\mathcal{S}$ during navigation, avoiding collisions with both known and unknown obstacles.
Two main strategies exist for designing safety-critical controllers in such cases. The first strategy treats unmeasured space as free space and focuses on avoiding previously known and newly detected obstacles. A common approach to this problem is the use of CBFs to formally guarantee collision avoidance. CBF-based quadratic programs (CBF-QPs) have emerged as a popular and computationally efficient method for enforcing safety-critical control objectives [2, 18]. However, a notable challenge emerges under limited sensing capabilities. Specifically, if an obstacle is detected in close proximity to the robot, particularly during rotational maneuvers, the CBF-QP may become infeasible as there might not be a feasible solution to avoid the obstacle due to the proximity and the robot’s dynamics.
@@
Assuming a safety-critical local controller $\pi$ interfacing with the global planner, we can now prove the properties of the resulting path from our proposed planner:
@@
Let $p^{\textup{ref}}:\mathcal{X}\to\mathcal{P}$ be the global reference path generated by the Visibility-Aware RRT* (Alg. 1), represented by a set of waypoints $\mathcal{P}=\{{\bm{x}}^{\textup{ref}}_{i}\}_{i=0}^{N}$. If there exists a local tracking controller $\pi$ with a maximum tracking error $\epsilon$ that can track the trajectory between consecutive waypoints, then the reference path $p^{\textup{ref}}$, when tracked by such controller $\pi$, is both collision-free and visibility-aware.
@@
The trajectory $\bm{\sigma}_{i}$ connecting consecutive waypoints ${\bm{x}}^{\textup{ref}}_{i}$ and ${\bm{x}}^{\textup{ref}}_{i+1}$ is generated by the LQR-CBF-Steer function. By construction, each trajectory satisfies both collision avoidance (20) and visibility (28) CBF constraints, considering the robot radius $r_{\text{r}}$ and the maximum tracking error $\epsilon$. Since (20) is satisfied, the trajectories remain within the set $\mathcal{C}_{\textup{col}}=\{{\bm{x}}\in\mathcal{X}:h_{\textup{col}}({\bm{x}})\geq 0\}$, where $h_{\textup{col}}$ is given by (19). This guarantees that the local controller $\pi$ follows the trajectory $\bm{\sigma}_{i}$ with a maximum tracking error of $\epsilon$ to remain collision-free (13). Moreover, since (28) is met, the local controller $\pi$ forces the robot to move along position trajectories $\bm{\sigma}_{i}$ from which the critical point (according to Definition 5) is visible before it is reached. This guarantees that the local controller $\pi$ can detect and avoid hidden obstacles that, in the worst-case scenario, are placed at the boundary of the local safe set $\mathcal{B}_{t}$. Consequently, the local controller $\pi$ can keep the robot within the known local collision-free sets $\mathcal{B}_{t}$ while tracking the reference path $p^{\textup{ref}}$, satisfying the visibility-aware condition (16). ∎
@@ VI RESULTS
We compare four planning methods in our experiments: (i) The A* algorithm with an obstacle distance cost [19] (Baseline 1), (ii) the standard LQR-RRT* algorithm [13] (Baseline 2), (iii) our proposed method without the visibility constraint (Baseline 3), and (iv) the propsed visibility-aware RRT*. To highlight the effect of the visibility constraint, we illustrate the reference paths generated by Baseline 3 and our method in Fig. 2. The results demonstrate that our method maintains fewer nodes in the tree compared to Baseline 3, as it only extends the nodes that satisfy the visibility constraint. This selective expansion leads to a more computationally efficient exploration of the configuration space [20]. [^2] https://github.com/tkkim-robot/visibility-rrt Furthermore, the final reference path generated by our method tends to take shallower turns around the known obstacles. This behavior complements the local tracking controller by providing a path that allows it to detect and avoid potential hidden obstacles.
To demonstrate the effectiveness of the generated paths when interfacing with the local tracking controllers, we employ two types of controllers to track the paths as discussed in Sec. V: a CBF-QP controller and the GateKeeper algorithm. We employ the dynamic unicycle model, which takes translational acceleration as input instead of velocity, in our tracking controllers to effectively capture the realistic dynamics and constraints of the robot [21].

### v2

@@ IV GLOBAL PATH PLANNER
In this section, we introduce the Visibility-Aware RRT* algorithm, which aims to solve the motion planning problem under limited sensing capabilities while satisfying the visibility condition given in (11). This algorithm builds upon the LQR-RRT* framework [13] by incorporating two CBFs that guide the search process to ensure satisfaction of both the collision-free (8) and visibility-aware (11) properties.
Following a common practice in motion planning problems, we use the (kinematic) unicycle model in the global path planner as a simple dynamical model, and the dynamic unicycle model, which takes translational acceleration as input instead of velocity, in the tracking controller to effectively capture more realistic dynamics and constraints of the robot.
@@ IV-A Visibility-Aware RRT*
The LQR-RRT* [13] is a sampling-based motion planning algorithm that incorporates the Linear Quadratic Regulator (LQR) controller as a steering function for nonlinear systems with non-holonomic constraints. Recently, Yang et al. [14] proposed an extension to the LQR-RRT* that incorporates CBF constraints during the planning process. Instead of formulating and solving Quadratic Programs (QPs) iteratively, as done in existing works [15], it checks whether the CBF constraints are satisfied and uses them as the termination condition for the steering process.
@@
for $k=1,\ldots,\texttt{{maxIter}}$ do
@@
  • Sample: Uniformly samples a random state ${\bm{x}}_{\text{rand}}=[x_{\text{rand}},y_{\text{rand}},\theta_{\text{rand}}]^{\top}$ within the configuration space to expand the tree towards.
  • ChooseParent (Alg. 2): Selects the best parent node ${\bm{x}}_{\text{min}}$ for the new node ${\bm{x}}_{\text{new}}$ from the set of nearby nodes $\mathcal{X}_{\text{near}}$ based on the cost and the feasibility of the connecting trajectory generated by the steering function. Given a trajectory $\bm{\sigma}=({\bm{x}}(t),{\bm{u}}(t))$, where $t\in[0,T]$ and $T$ is the duration of the trajectory, the evaluated quadratic cost is defined as:
    $\texttt{cost}(\bm{\sigma})=\int_{0}^{T}{\bm{x}}(t)^{\top}Q{\bm{x}}(t)+{\bm{u}}(t)^{\top}R{\bm{u}}(t)\,\mathrm{d}t,$ (12)
    where $Q\succeq 0$ and $R\succ 0$ are weight matrices.
  • Rewire (Alg. 3): Performs the rewiring process by considering the new node ${\bm{x}}_{\text{new}}$ as a potential parent for the nearby nodes in $\mathcal{X}_{\text{near}}$. If a trajectory from ${\bm{x}}_{\text{new}}$ to any nearby node ${\bm{x}}_{\text{near}}\in\mathcal{X}_{\text{near}}$ is found with a smaller cost (12), the tree is optimized by rewiring ${\bm{x}}_{\text{near}}$ to ${\bm{x}}_{\text{new}}$.
@@
The LQR-CBF-Steer is the key function that allows the algorithm to use LQR and CBF (2) as an extension heuristic between two nodes ${\bm{x}}_{\text{start}}$ and ${\bm{x}}_{\text{next}}$ (see Alg. 4). It first linearizes the dynamics around ${\bm{x}}_{\text{next}}$ and computes the optimal gain $K_{\text{lqr}}$ by solving the algebraic Riccatti equation. The function then employs the LQR controller $\pi_{\text{lqr}}$ to generate a sequence of intermediate states $\mathcal{X}^{\prime}=\{{\bm{x}}_{\text{start}},{\bm{x}}^{\prime}_{1},\ldots,{\bm{x}}^{\prime}_{T}\}$ and control inputs $\mathcal{U}^{\prime}=\{{\bm{u}}^{\prime}_{1},\ldots,{\bm{u}}^{\prime}_{T}\}$, steering the state from ${\bm{x}}_{\text{start}}$ towards ${\bm{x}}_{\text{next}}$ based on the linearized dynamics. Throughout this process, it verifies that all CBF constraints are satisfied for each state-action pair $({\bm{x}}^{\prime}_{t},{\bm{u}}^{\prime}_{t})$ at each discretized time step $t$. If any constraint is violated, the steering extension is immediately terminated [14].
@@ IV-B HOCBF for Collision Avoidance
The unicycle model is as follows:
@@
Hereafter, we abuse the notation and use ${\bm{x}}_{t}$ and ${\bm{u}}_{t}$ to refer to ${\bm{x}}^{\prime}_{t}$ and ${\bm{u}}^{\prime}_{t}$ in the context of the LQR-CBF-Steer function. Let one of the obstacles with radius $l_{\text{obs}}\in\mathbb{R}^{+}$ be located at $(x_{\text{obs}},y_{\text{obs}})\in\mathbb{R}^{2}$. The constraint function (3a) for collision avoidance, considering the robot radius $l_{\text{robot}}$ and the maximum tracking error $\epsilon$ mentioned in Sec. III, is:
$h_{\textup{col}}({\bm{x}}_{t})=(x_{t}-x_{\text{obs}})^{2}+(y_{t}-y_{\text{obs}})^{2}-(l_{\text{obs}}+l_{\text{robot}}+\epsilon)^{2}.$ (14)
Given the dynamics (13), the function $h_{\textup{col}}({\bm{x}}_{t})$ in (14) has relative degree $r=1$ for the control input $v$ and $r=2$ for the control input $\omega$. To simplify the construction of the CBF constraint, we fix the translational velocity $v$ to a constant value and only control the angular velocity $\omega$ [15].
Then, the HOCBF constraint (5) with $r=2$ is [15]:
@@
$+\gamma_{1}L_{f}h_{\textup{col}}({\bm{x}}_{t})+\gamma_{2}h_{\textup{col}}({\bm{x}}_{t})\geq 0,$ (15)
where $\gamma_{1}$ and $\gamma_{2}$ are positive constants that are designed to satisfy the HOCBF condition (5), and
@@ IV-C1 Time-to-Reach
Given the robot’s current position ${\bm{r}}_{t}=[x_{t},y_{t}]^{\top}$ and the critical point ${\bm{x}}_{c}$, let $\Delta{\bm{r}}=\sqrt{(x_{t}-x_{c})^{2}+(y_{t}-y_{c})^{2}}$ denote the Euclidean distance between them. To account conservatively for the robot radius $l_{\text{robot}}$ and the maximum tracking error $\epsilon$ of the local controller $\pi$, we define $\Delta d=\Delta{\bm{r}}-l_{\text{robot}}-\epsilon$. The time-to-reach is:
@@ IV-C2 Time-to-Rotate
where $\bar{\omega}_{t}$ is the average angular velocity computed by simulating the LQR controller $\pi_{\text{lqr}}$ for a rotation of $\Delta\theta$ starting from time $t$, i.e., $\bar{\omega}_{t}=\frac{1}{T_{\omega}}\sum_{\tau=t}^{t+T_{\omega}}\omega_{\tau}^{\text{lqr}}$. $\omega_{\tau}^{\text{lqr}}$ denotes the angular velocity generated by the LQR controller $\pi_{\text{lqr}}$ at time step $\tau$, and $T_{\omega}$ denotes the total number of time steps required for the rotation.
@@ IV-C3 Visibility CBF
Then, the CBF constraint with $r=1$ (2) is defined as:
$\psi_{\textup{vis}}({\bm{x}}_{t})=\dot{h}_{\textup{vis}}({\bm{x}}_{t})+\gamma_{3}\,h_{\textup{vis}}({\bm{x}}_{t})\geq 0,$ (22)
where $\gamma_{3}$ is positive constant that is designed to satisfy the CBF condition. The gradient $\dot{h}_{\textup{vis}}({\bm{x}}_{t})$ can be derived as:
@@
Proof.
Let us verify that condition (2) holds for all ${\bm{x}}\in\mathcal{X}$. Note that $L_{f}h_{\textup{vis}}({\bm{x}}_{t})$ is bounded in $[-1,1]$, and $L_{g}h_{\textup{vis}}({\bm{x}}_{t})$ takes the value ±$\pm 1/\bar{\omega}_{t}$ depending on $\theta_{t}$ and $\theta_{c}$. Plus, $\bar{\omega}_{t}$ is bounded because it is computed from the average LQR control input, which converges to zero over time, and the largest control input is bounded in the control space $\mathcal{X}$. Therefore, a sufficiently large control bound on the angular velocity $\omega_{\max}$ satisfying $\omega_{\max}\geq\bar{\omega}_{t}$ guarantees that there exists a control input $\omega_{t}$ such that the condition (2) is satisfied, even at the boundary $h_{\textup{vis}}({\bm{x}}_{t})=0$. Hence, $h_{\textup{vis}}$ is a valid CBF. ∎
@@ V LOCAL TRACKING CONTROLLER
After the global planning cycle described in Sec. IV, the controller $\pi$ tracks the waypoints ${\bm{x}}_{j}^{\text{ref}}\in\mathcal{P}$ enabling the robot to navigate towards the goal. During navigation, the robot utilizes its onboard sensor to create a map of the environment and detect obstacles on-the-fly. The environment may contain unknown obstacles that were not considered during the global planning phase. The controller must ensure that the robot remains within the true collision-free set $\mathcal{S}$ during navigation, avoiding collisions with both known and unknown obstacles.
Two main strategies exist for designing safety-critical tracking controllers in such cases. The first strategy treats unmeasured space as free space and focuses on avoiding previously known and newly detected obstacles. A common approach to this problem is the use of CBFs to formally guarantee collision avoidance. CBF-based quadratic programs (CBF-QPs) have emerged as a popular and computationally efficient method for enforcing safety-critical control objectives [2]. However, a notable challenge emerges under limited sensing capabilities. Specifically, if an obstacle is detected in close proximity to the robot, particularly during rotational maneuvers, the CBF-QP may become infeasible as there might not be a feasible solution to avoid the obstacle due to the proximity and the robot’s dynamics.
@@
Assuming a safety-critical tracking controller $\pi$ interfacing with the global planner, we can now prove the properties of the resulting path from our proposed planner:
@@
Let $p^{\textup{ref}}$ be the global reference path generated by the Visibility-Aware RRT* (Alg. 1). If there exists a local tracking controller $\pi$ that can track the trajectory between consecutive waypoints with a maximum tracking error $\epsilon$, then the reference path $p^{\textup{ref}}$, when tracked by such controller $\pi$, is both collision-free and visibility-aware.
@@
By construction, each trajectory $\bm{\sigma}_{j}$ generated by the LQR-CBF-Steer function satisfies both collision avoidance (15) and visibility (22) CBF constraints, considering the robot radius $l_{\text{robot}}$ and the maximum tracking error $\epsilon$. Since (15) is satisfied, the trajectories remain within the set $\mathcal{C}_{\textup{col}}$. This guarantees that the controller $\pi$ follows the trajectory $\bm{\sigma}_{j}$ with a maximum tracking error of $\epsilon$ to remain collision-free (8). Moreover, since (22) is met, the controller $\pi$ forces the robot to move along position trajectories $\bm{\sigma}_{j}$ from which the critical point (according to Definition 5) is visible before it is reached. This guarantees that the tracking controller $\pi$ can detect and avoid hidden obstacles that, in the worst-case scenario, are placed at the boundary of the local safe set $\mathcal{B}_{t}$. Consequently, the controller $\pi$ can keep the robot within the known local collision-free sets $\mathcal{B}_{t}$ while tracking the reference path $p^{\textup{ref}}$, satisfying the visibility-aware condition (11). ∎
@@ VI RESULTS
We compare four planning methods in our experiments: (i) The A* algorithm with an obstacle distance cost [17] (Baseline 1), (ii) the standard LQR-RRT* algorithm [13] (Baseline 2), (iii) our proposed method without the visibility constraint (Baseline 3), and (iv) the propsed Visibility-Aware RRT*. To highlight the effect of the visibility constraint, we illustrate the reference paths generated by Baseline 3 and our method in Fig. 3. The results demonstrate that our method maintains fewer nodes in the tree compared to Baseline 3, as it only extends the nodes that satisfy the visibility constraint. This selective expansion leads to a more computationally efficient exploration of the configuration space. Furthermore, the final reference path generated by our method tends to take shallower turns around the known obstacles. This behavior complements the local tracking controller by providing a path that allows it to detect and avoid potential hidden obstacles.
To demonstrate the effectiveness of the generated paths when interfacing with the local tracking controllers, we employ two types of controllers to track the paths as discussed in Sec. V: a CBF-QP controller and the Gatekeeper algorithm. We adopt the dynamic unicycle model for our controllers.
@@
In addition to simulations, we conducted hardware experiments using a ground rover modeled as a dynamic unicycle (see Fig. 1). An RGB-D camera was used for localization and obstacle detection with a small safety buffer to mitigate sensor noise. All navigation tasks were operated onboard without a motion capture system. Hidden obstacles were not detected in time while tracking paths that were generated by baseline global path planners, leading to collisions. In Fig. 1, the naive path was generated by Baseline 3. In contrast, the paths generated by our method were able to avoid collisions by accounting for perception limitations. The detailed system setup and results are available in the supplementary video.